solve the equation.4x-6=5(x-5)
Answer by Cromlix(4381)

4x-6=5(x-5)
4x - 6 = 5x - 25
Collect like terms
4x - 5x = -25 + 6
-x = - 19
x = 19
hope this helps
:-)